# On the Parameter Identifiability of Partially Observed Linear Causal Models

### v1

@@ 4.2 Parameterization Trick of Covariance Matrix
In this section, we will introduce how trek rules can be employed to parameterize $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ while the unit variance constraint on latent variables can be elegantly satisfied. We start with the definition of trek. For readers who are not familiar with treks, please refer to Appendix B.1 for an illustrative example.
@@
At this point, we are able to parameterize each entry of $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ using ($F,\{\sigma_{ii}\}_{i=1}^{n+m}$), instead of ($F,\Omega$), by making use of the (simple) trek rule [49], as follows:
@@
where $\mathcal{S}(\mathsf{V_{i}},\mathsf{V_{j}})$ is the set of all simple treks between $\mathsf{V_{i}}$ and $\mathsf{V_{j}}$, and $f^{P}$ is the path monomial along $P$ defined as $f^{P}:=\Pi_{k\rightarrow l\in P}f_{kl}$. By this form of parameterization, we can simply set all entries of $\{\sigma_{ii}\}_{i=1}^{n+m}$ as 1 (which is equivalent to requiring all variables to have unit variance), such that the constraint in Equation 3 can be automatically satisfied. For a better understanding of how to use the simple trek rule for parameterization, we provide an example as follows.
@@
As shown in Example 6, entries of $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ can be written as a function of entries of $F$, by assuming all variables to have unit variance. This frees us from adding penalty terms to optimization and we will show later in Section 5 that it does achieve better empirical results.
@@ 5.1 Setting and Evaluation Metric
Table 1: Experimental result on synthetic data. We note that, to the best of our knowledge, there is no existing method that can achieve the same goal as ours. Thus, Estimator-LM serves as the baseline.
@@
Method Estimator-LM Estimator-TR
2k 0.030 (0.020) 0.007 (0.009)
GS Case 5k 0.030 (0.010) 0.002 (0.002)
10k 0.020 (0.010) 0.001 (0.001)
(a) MSE up to group sign (mean (std)).
MSE up to orthogonal
Method Estimator-LM Estimator-TR
2k 0.050 (0.010) 0.021 (0.010)
OT Case 5k 0.040 (0.010) 0.018 (0.010)
10k 0.040 (0.010) 0.013 (0.010)
(b) MSE up to orthogonal transformation (mean (std)).
Figure 4: Estimated edge coefficients by the proposed method for Big Five human personality dataset.
As the optimization in Eq 3 is nonconvex, we will rely on 30 random starts and choose the one with the best likelihood. We report the performance of the proposed method with two different ways to constrain latent variables to have unit variance. (i) Parameter Estimator with Lagrange multiplier, referred to as Estimator-LM, and (ii) Parameter Estimator with Trek Rule, referred to as Estimator-TR.
It is worth noting that our setting is very general in that we allow latent variables and observed variables to be causally connected in a very flexible way, and we consider the identification of parameters of edges that can involve both observed and latent variables. Therefore, to the best of our knowledge, no current method can achieve the same goal. As such, we mainly focus on comparing our estimation result with the ground truth parameters and also use our Estimator-LM to serve as a baseline for comparison. This also shows the novelty of the proposed method.
We use two MSE-based metrics defined as follows. MSE upto group sign: suppose the ground truth parameter is $F$ and our estimation is $\hat{F}$, the MSE upto group sign is defined as $(\||F|-|\hat{F}|\|^{2}_{2})/\|F\|_{0}$, where ||$\cdot$|| ⋅ | takes element wise absolute value, $\|\cdot\|_{2}$ denotes the Frobenius norm and $\|\cdot\|_{0}$ denotes the number of nonzero entries of a matrix. MSE upto orthogonal transformation: the MSE upto orthogonal transformation is defined as $\min_{Q:Q^{T}Q=I}(\||F|-|Q\hat{F}|\|^{2}_{2})/\|F\|_{0}$.
@@ 5.2 Synthetic Data Performance
We report the performance using synthetic data in Tab 1(a) and 1(b), where our Estimator-TR achieves very good identification performance. For example, in the GS scenario with 10k samples, our Estimator-TR achieves 0.001 MSE up to group sign. In contrast, the baseline Estimator-LM is less satisfying; it achieves 0.020 MSE up to group sign using 10k samples. The good performance by Estimator-TR not only validates our estimation method, but also empirically verifies our identifiability theory. Plus, the performance gap between Estimator-LM and Estimator-TR illustrates the necessity of the parameterization trick for the covariance matrix.
@@ 5.3 Misspecification Behavior
As for violation of normality, we use uniform noise terms for the underlying model, and thus the distribution is not jointly gaussian anymore. We aim to see to what extent can the proposed method still recover the correct parameters. The result is shown in Table 2 in the Appendix, which shows even when the normality is violated, we can still estimate the parameters quite well.
To simulate the violation of linearity, we employ the leaky ReLU function during the generation process, as $\mathsf{V_{i}}=\text{LRELU}(\sum\nolimits_{\mathsf{V_{j}}\in\text{Pa}(\mathsf{V_{i}})}f_{ji}\mathsf{V_{j}}+\epsilon_{\mathsf{V_{i}}})$, $\text{LRELU}(x)=\max(\alpha x,x)$. When $\alpha$ is close to 1, the function is close to a linear one, and when $\alpha$ is close to 0, the model is very nonlinear. The result is shown in Table 3 and we found that our estimation method is quite robust to small violations of linearity. For example, in GS case with 10k sample size, if we set $\alpha=0.8$, we still get a small MSE of 0.003. Even when $\alpha$ decreases to 0.6, the MSE is around 0.009, which is still small. However, when $\alpha$ is decreased to $0.3$, the model is considerably nonlinear, and the MSE increases to 0.040.
@@ 5.4 Implementation Details, Runtime Analysis, and Scalability
Our code is based on Python3.7 and PyTorch [35]. Data is standardized and the optimization in Eq 3 and MSE calculation (up to orthogonal transformation) are solved by Adam [26], with a learning rate of $0.02$. We conduct all the experiments with single Intel(R) Xeon(R) CPU E5-2470. All experiments can be finished within $2$ hours. We note that our method is very computationally efficient. First, the computational cost is almost irrelevant to sample size: we only need to calculate the sample covariance matrix once and cache it for further use during the optimization. Plus, our estimation method can handle a large number of variables. For example, the running time of our method are roughly 10 seconds, 2 minutes, and 10 minutes for 20 variables, 50 variables, and 100 variables respectively. For 300 variables, which is a considerably large number for typical experiments considered in causal discovery papers, the estimation can still be finished within around one hour.
It is also worth noting that model misspecifications do not influence the computation cost of our method. We briefly discuss the efficiency of checking whether conditions in Theorem 1 hold, together with what if conditions do not hold in solving real-life problems in Appendix A.6 and A.7.
@@ 5.5 Real-World Data Performance
In this section, we employ a famous psychometric dataset - Big Five dataset https://openpsychometrics.org/, to validate our method. It consists of 50 indicators and close to 20,000 data points. There are five dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (O-C-E-A-N). Each is measured with 10 indicators. Data is standardized.We employ the RLCD method [17] to determine the MEC and GIN [56] to decide the remaining directions. Then we employ the proposed Estimator-TR to estimate all the edge coefficients.
@@ A.1 Proof of Lemma 1
Proof of Lemma 1.
We adopt the definition of matrices $M$ and $N$ in Proposition 1. Let $\Lambda$ be a diagonal matrix with diagonal entries. We rescale the matrices in the following way: $\hat{A}\coloneqq\Lambda^{-1}A\Lambda$, $\hat{B}\coloneqq\Lambda^{-1}B$, $\hat{C}\coloneqq C\Lambda$, $\hat{D}\coloneqq D$, $\hat{\Omega}_{\mathbf{L}}\coloneqq\Lambda^{2}\Omega_{\mathbf{L}}=\Lambda\Omega_{\mathbf{L}}\Lambda$, and $\hat{\Omega}_{\mathbf{X}}\coloneqq\Omega_{\mathbf{X}}$. We then have
$\hat{M}$ $=\left(I-\hat{A}-\hat{B}(I-\hat{D})^{-1}\hat{C}\right)^{-1}$
$=\left(\Lambda^{-1}\Lambda-\Lambda^{-1}AQ-(\Lambda^{-1}B)(I-D)^{-1}(C\Lambda)\right)^{-1}$
@@
$\hat{N}$ $=\left((I-\hat{A})\hat{C}^{-1}(I-\hat{D})-\hat{B}\right)^{-1}$
@@
By Proposition 1, the latent covariance matrix $\hat{\Sigma}_{\mathbf{L}}$ after rescaling of the parameters is given by
$\hat{\Sigma}_{\mathbf{L}}$ $=\hat{M}^{T}\hat{\Omega}_{\mathbf{L}}\hat{M}+\hat{N}^{T}\hat{\Omega}_{\mathbf{X}}\hat{N}$
$=(\Lambda^{T}M^{T}\Lambda^{-T})(\Lambda\Omega_{\mathbf{L}}\Lambda)(\Lambda^{-1}M\Lambda)+\Lambda^{T}N^{T}\Omega_{\mathbf{X}}N\Lambda$
$=\Lambda(M^{T}\Omega_{\mathbf{L}}M+N^{T}\Omega_{\mathbf{X}}N)\Lambda$
$=\Lambda\Sigma_{L}\Lambda.$
This implies that the variance of each latent variable $\mathsf{L}_{i}$ is scaled by $\Lambda_{ii}^{2}$. By Proposition 1, the observed covariance matrix $\hat{\Sigma}_{\mathbf{X}}$ after rescaling of the parameters is given by
$\hat{\Sigma}_{\mathbf{X}}$ $=(I-\hat{D})^{-T}\Big{(}\hat{\Omega}_{\mathbf{X}}+\hat{B}^{T}\hat{\Sigma}_{L}B+\hat{\Omega}_{\mathbf{X}}\hat{N}\hat{B}+\hat{B}^{T}\hat{N}^{T}\hat{\Omega}_{\mathbf{X}}\Big{)}(I-\hat{D})^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+(\Lambda^{-1}B)^{T}(\Lambda\Sigma_{L}\Lambda)(\Lambda^{-1}B)+\Omega_{\mathbf{X}}(N\Lambda)(\Lambda^{-1}B)+(\Lambda^{-1}B)^{T}(N\Lambda)^{T}\Omega_{\mathbf{X}}\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+B^{T}\Sigma_{L}B+\Omega_{\mathbf{X}}NB+B^{T}N^{T}\Omega_{\mathbf{X}}\Big{)}(I-D)^{-1}$
@@
This indicates that the rescaled parameters can still entailed the observations $\Sigma_{\mathbf{X}}$. Therefore, we can rescale the variance of any latent variables and change the coefficients in specific ways such that the entailed observation $\Sigma_{\mathbf{X}}$ is unchanged, indicating that the variance of latent variables can never be identified. ∎
@@ A.2 Proof of Theorem 1
The structure identifiability part is that if $\mathcal{G}$ satisfies Condition 1 and Condition 2, the structure of $\mathcal{G}$ can be identified upto the Markov equivalence class of $\mathcal{O}_{\text{min}}(\mathcal{O}_{s}(\mathcal{G}))$, which is by Theorem 12 in [17].
Next we will focus on the proof of parameter identifiability part, i.e., for any DAG in the equivalence class, if (i) and (ii) in Theorem 1 are satisfied, the parameters are identifiable (up to group sign). Without loss of generality, we assume all variables have unit variance and zero mean.

### v2

@@ 4.2 Parameterization Trick of Covariance Matrix
In this section, we introduce how trek rules can be employed to parameterize $\Sigma_{\mathbf{X}}$ while the unit variance constraint on latent variables in Eq. (5) can be elegantly satisfied. We start with the definition of trek. For readers who are less familiar with treks, please refer to Appendix B.1 for examples.
@@
At this point, we are able to parameterize each entry of $\Sigma_{\mathbf{X}}$ using ($F,\{\sigma_{ii}\}_{i=1}^{n+m}$), instead of ($F,\Omega$), by making use of the (simple) trek rule [50], as follows:
@@
where $\mathcal{S}(\mathsf{V_{i}},\mathsf{V_{j}})$ is the set of all simple treks between $\mathsf{V_{i}}$ and $\mathsf{V_{j}}$, and $f^{P}$ is the path monomial along $P$ defined as $f^{P}:=\Pi_{k\rightarrow l\in P}f_{kl}$.
By this form of parameterization, we can simply set all entries of $\{\sigma_{ii}\}_{i=1}^{n+m}$ as 1 (which is equivalent to requiring all variables to have unit variance), such that the constraint in Eq. (5) can be automatically satisfied. For a better understanding of how to use the simple trek rule for parameterization, we provide an example as follows.
@@ 5.1 Setting and Evaluation Metric
Table 1: Experimental result on synthetic data using MSE (mean (std)).
@@
Method Estimator Estimator-TR
2k 0.0023 (0.002) 0.0012 (0.0005)
GS Case 5k 0.0014 (0.002) 0.0005 (0.0005)
10k 0.0012 (0.001) 0.0003 (0.0004)
(a) MSE up to group sign indeterminacy.
MSE up to orthogonal.
Method Estimator Estimator-TR
2k 0.0278 (0.008) 0.0355 (0.015)
OT Case 5k 0.0194 (0.002) 0.0352 (0.012)
10k 0.0182 (0.003) 0.0351 (0.015)
(b) MSE up to orthogonal transformation.
Figure 4: Estimated edge coefficients by the proposed method for Big Five human personality dataset. Variables whose name starts with "L" are latent variables while the others are observed variables.
As the optimization in Eq (4) is nonconvex, we will rely on 30 random starts and choose the one with the best likelihood. We report the performance of the proposed method with two different objectives. (i) Parameter Estimator with objective defined in Eq. (4), referred to as Estimator, and (ii) Parameter Estimator with objective defined in Eq. (5) and Trek Rule parameterization trick in Eq (6), referred to as Estimator-TR.
It is worth noting that our setting is very general in that we allow latent variables and observed variables to be causally connected in a very flexible way, and we consider the identification of parameters of edges that can involve both observed and latent variables. Therefore, to the best of our knowledge, no current method can achieve the same goal to serve as the baseline (which also shows the novelty of the proposed method). As such, we mainly focus on comparing our estimation result with the ground truth parameters. We use two MSE-based metrics defined as follows.
MSE up to group sign: suppose the ground truth parameter is $F$ and our estimation is $\hat{F}$. The MSE up to group sign is defined as $\frac{\||F|-|\hat{F}|\|^{2}_{2}}{\|F\|_{0}},$ where ||$\cdot$|| ⋅ | takes element wise absolute value, $\|\cdot\|_{2}$ denotes the Frobenius norm and $\|\cdot\|_{0}$ denotes the number of nonzero entries of a matrix.
MSE up to orthogonal transformation: the MSE up to orthogonal transformation is defined as
$\min_{Q:Q^{T}Q=I}~{}\frac{\||F_{\mathbf{L}\mathbf{L}}|-|Q^{T}\hat{F}_{\mathbf{L}\mathbf{L}}Q|\|^{2}_{2}+\||F_{\mathbf{L}\mathbf{X}}|-|Q^{T}\hat{F}_{\mathbf{L}\mathbf{X}}|\|^{2}_{2}+\||F_{\mathbf{X}\mathbf{L}}|-|\hat{F}_{\mathbf{X}\mathbf{L}}Q|\|^{2}_{2}+\||F_{\mathbf{X}\mathbf{X}}|-|\hat{F}_{\mathbf{X}\mathbf{X}}|\|^{2}_{2}}{\|F\|_{0}},$ (7)
where the optimization is solved by Adam [27] and the orthogonal matrix $Q$ can be directly parameterized in PyTorch.
@@ 5.2 Synthetic Data Performance
We report the performance using synthetic data in Tables 1(a) and 1(b), where both our Estimator and Estimator-TR achieve very good identification performance. For example, in the GS scenario with 10k samples, our Estimator achieves 0.0.0012 MSE up to group sign and our Estimator-TR achieves 0.0003 MSE up to group sign. The good performance by Estimator and Estimator-TR not only validates our estimation method, but also empirically verifies our identifiability theory.
@@ 5.3 Misspecification Behavior
As for violation of normality, we use uniform noise terms for the underlying model, and thus the distribution is not jointly Gaussian anymore. We aim to see to what extent can the proposed method still recover the correct parameters. The result is shown in Table 2 in the Appendix, which shows even when the normality is violated, we can still estimate the parameters pretty well. The reason lies in that our proposed asymptotic identifiability result holds true, even when we do not assume Gaussianity; as we only make use of the second-order statistics of the distribution, the additive noise in Definition 1 can follow any other continuous distribution.
To simulate the violation of linearity, we employ the leaky ReLU function during the generation process, as $\mathsf{V_{i}}=\text{LRELU}(\sum\nolimits_{\mathsf{V_{j}}\in\text{Pa}(\mathsf{V_{i}})}f_{ji}\mathsf{V_{j}}+\epsilon_{\mathsf{V_{i}}})$, $\text{LRELU}(x)=\max(\alpha x,x)$. When $\alpha$ is close to 1, the function is close to a linear one, and when $\alpha$ is close to 0, the model is very nonlinear. The result is shown in Table 3 and we found that our estimation method is quite robust to small violations of linearity. For example, for Estimator-TR in GS case with 10k sample size, if we set $\alpha=0.8$, we still get a small MSE of 0.001. Even when $\alpha$ decreases to 0.6, the MSE is around 0.005, which is still small. However, when $\alpha$ is decreased to $0.3$, the underlying model is considerably nonlinear, and the MSE increases to 0.027.
@@ 5.4 Implementation Details, Runtime Analysis, and Scalability
Our code is based on Python3.7 and PyTorch [37]. Data is standardized and the optimization in Eqs. (4), (5), and (7) are solved by Adam [27], with a learning rate of $0.02$. We conduct all the experiments with single Intel(R) Xeon(R) CPU E5-2470. All experiments can be finished within $2$ hours. We note that our method is very computationally efficient. First, the computational cost is almost irrelevant to sample size: we only need to calculate the sample covariance matrix once and cache it for further use during the optimization. Plus, our estimation method can handle a large number of variables. For example, the running time of our method are roughly 10 seconds, 2 minutes, and 10 minutes for 20 variables, 50 variables, and 100 variables respectively. For 300 variables, which is a considerably large number for typical experiments considered in causal discovery papers, the estimation can still be finished within around one hour.
It is also worth noting that model misspecifications do not influence the computation cost of our method. We briefly discuss the efficiency of checking whether conditions in Theorem 3 hold, together with what if conditions do not hold in solving real-life problems in Appendices A.8 and A.9.
@@ 5.5 Real-World Data Performance
In this section, we employ a famous psychometric dataset - Big Five dataset https://openpsychometrics.org/, to validate our method. It consists of 50 indicators and close to 20,000 data points. There are five dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (O-C-E-A-N). Each is measured with 10 indicators. Data is standardized.We employ the RLCD method [18] to determine the MEC and GIN [58] to decide the remaining directions. Then we employ the proposed Estimator-TR to estimate all the edge coefficients. The structure satisfies the condition in Theorem 5 so we know that the parameters are identifiable.
@@ 6 Conclusion
7 Acknowledgement
This material is based upon work supported by NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Salesforce, Apple Inc., Quris AI, Florin Court Capital, and the MBZUAI-WIS grant.
@@ A.1 Proof of Theorem 1
Proof of Theorem 1.
[…]
$\tilde{M}$ $=\left(I-\tilde{A}-\tilde{B}(I-\tilde{D})^{-1}\tilde{C}\right)^{-1}$
$=\left(\Lambda^{-1}\Lambda-\Lambda^{-1}A\Lambda-(\Lambda^{-1}B)(I-D)^{-1}(C\Lambda)\right)^{-1}$
@@
$\tilde{N}$ $=\left((I-\tilde{A})\tilde{C}^{-1}(I-\tilde{D})-\tilde{B}\right)^{-1}$
@@
By Proposition 1, the latent covariance matrix $\tilde{\Sigma}_{\mathbf{L}}$ after rescaling of the parameters is given by
$\tilde{\Sigma}_{\mathbf{L}}$ $=\tilde{M}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}\tilde{M}+\tilde{N}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\tilde{N}$
$=(\Lambda^{T}M^{T}\Lambda^{-T})(\Lambda\Omega_{\epsilon_{\mathbf{L}}}\Lambda)(\Lambda^{-1}M\Lambda)+\Lambda^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}N\Lambda$
$=\Lambda(M^{T}\Omega_{\epsilon_{\mathbf{L}}}M+N^{T}\Omega_{\epsilon_{\mathbf{X}}}N)\Lambda$
$=\Lambda\Sigma_{\mathbf{L}}\Lambda.$
This implies that the variance of each latent variable $\mathsf{L}_{i}$ is scaled by $\Lambda_{ii}^{2}$. By Proposition 1, the observed covariance matrix $\tilde{\Sigma}_{\mathbf{X}}$ after rescaling of the parameters is given by
$\tilde{\Sigma}_{\mathbf{X}}$ $=(I-\tilde{D})^{-T}\Big{(}\tilde{B}^{T}\tilde{\Sigma}_{\mathbf{L}}B+\tilde{\Omega}_{\epsilon_{\mathbf{X}}}+\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\tilde{N}\tilde{B}+\tilde{B}^{T}\tilde{N}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\Big{)}(I-\tilde{D})^{-1}$
$=(I-D)^{-T}\Big{(}(\Lambda^{-1}B)^{T}(\Lambda\Sigma_{\mathbf{L}}\Lambda)(\Lambda^{-1}B)$
$\qquad\qquad\qquad\qquad+\Omega_{\epsilon_{\mathbf{X}}}+\Omega_{\epsilon_{\mathbf{X}}}(N\Lambda)(\Lambda^{-1}B)+(\Lambda^{-1}B)^{T}(N\Lambda)^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}B^{T}\Sigma_{\mathbf{L}}B+\Omega_{\epsilon_{\mathbf{X}}}+\Omega_{\epsilon_{\mathbf{X}}}NB+B^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-D)^{-1}$
@@ A.2 Proof of Theorem 2
See 2
Proof of Theorem 2.
[…]
$\tilde{A}\coloneqq S^{-1}AS,\quad\tilde{B}\coloneqq S^{-1}B,\quad\tilde{C}\coloneqq CS,\quad\tilde{D}\coloneqq D,\quad\tilde{\Omega}_{\epsilon_{\mathbf{L}}}\coloneqq S^{2}\Omega_{\epsilon_{\mathbf{L}}},\quad\text{and}\quad\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\coloneqq\Omega_{\epsilon_{\mathbf{X}}}.$
Note that $S$ is an invertible diagonal matrix. By Theorem 1, we have $\tilde{\Sigma}_{\mathbf{X}_{\mathcal{G}}}=\Sigma_{\mathbf{X}_{\mathcal{G}}}$ and $\tilde{\Sigma}_{\mathbf{L}_{\mathcal{G}}}=S\Sigma_{\mathbf{L}_{\mathcal{G}}}S$, and thus $(\tilde{\Sigma}_{\mathbf{L}_{\mathcal{G}}})_{ii}=(\Sigma_{\mathbf{L}_{\mathcal{G}}})_{ii},~{}\forall i\in[m]$. ∎
A.3 Proof of Theorem 3
The structure identifiability part is that if $\mathcal{G}$ satisfies Condition 1 and Condition 2, the structure of $\mathcal{G}$ can be identified up to the Markov equivalence class of $\mathcal{O}_{\text{a}}(\mathcal{O}_{s}(\mathcal{G}))$, which is by Theorem 12 in [18].
Next we will focus on the proof of parameter identifiability part, i.e., for any DAG in the equivalence class, if (i) and (ii) in Theorem 3 are satisfied, the parameters are identifiable (up to group sign). Without loss of generality, we assume that all variables have unit variance and zero mean. The reason is that if we can show that the parameters are identfiable (up to group sign) under this assumption, then it is straightforward to show that they are also identifiable under the original assumption where $\Omega_{\epsilon_{\mathbf{L}}}=I$.